How do I solve log2(x+5) -3 = log2(x-2) ?
Thanks for your help

Answer by Alan3354(21583)

How do I solve log2(x+5) -3 = log2(x-2)
--------------------
Subtracting 3 from a log is dividing the antilog by the base to the power of 3, that's the secret.
log2(x+5) -3 = log2(x-2)
log2((x+5)/8) = log2(x-2)
So, (x+5)/8 = x-2
x+5 = 8x-16
21 = 7x
x = 3
-----------
Check:
log2(8) - 3 =? log2(1)
3-3 = 0, so it's correct.